OpenStudy (anonymous):

Find the constant of variation k for the inverse variation. Then write an equation for the inverse variation. Y = 4.5 when x = 3 A) k = 1.5; y = 1.5/x B) k = 13.5; xy – 13.5 C) k = 1.5; y = 1.5x D) k = 13.5; 13.5y = x

Nnesha (nnesha):

okay so do you know what equation u have to use ??

Nnesha (nnesha):

hmm that is slope intercept form equation

Nnesha (nnesha):

\[\huge\color{blue}{{\rm y = }\rm \frac{ k }{ x }}\] this is equation for inverse variation u have to find k(constant of variation ) so replace y and x by the value that they gave you then solve for k

OpenStudy (anonymous):

C

Nnesha (nnesha):

http://openstudy.com/code-of-conduct @RogerYE welcome to open study!!!

Nnesha (nnesha):

let me know if you didn't understand @Answers101

OpenStudy (anonymous):

I don't understand @Nnesha

Nnesha (nnesha):

okay i gave u the equation right and u have y and x values u need to find k replace y by 4.5 and x by 3 then solve for k

OpenStudy (anonymous):

I think the answer is either B or D because 13.5 divided by 3 is 4.5

Nnesha (nnesha):

what is k value ??

OpenStudy (anonymous):

13.5

Nnesha (nnesha):

yes right now replace k by 13.5 in the equation that i gave u

OpenStudy (anonymous):

Y = k/x So 4.5 = 13.5/3 right?

Nnesha (nnesha):

so which one ur answer ??

OpenStudy (anonymous):

D ?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

look at ur answer choices there is mistake somewhere check it

OpenStudy (anonymous):

B

Nnesha (nnesha):

nope i'm saying there is mistake somewhere look at again there is typo or error

OpenStudy (anonymous):

On B I meant = instead of -

Nnesha (nnesha):

yess so its  is it right or not ??

OpenStudy (anonymous):

Yes